OpenStudy (anonymous):

Match the sequence (term) with the correct type of sequence (definition). Match Term Definition 2, 4, 6, 8, ... A) Geometric, common ratio is 2 28, 24, 20, 16, ... B) Geometric, common ratio is 0.5 2, 4, 8, 16, ... C) Arithmetic, common difference is -4 48, 24, 12, 6, ... D) Arithmetic, common difference is 2

OpenStudy (arbershabani97):

1st part is arithmetic because you only add 2 2nd is arithmetic because you only add -4 3rd is geometric because you are multiplying each one by 2 (2*2=4)(4*2=8) 4th is geometric because you are multiplying each with 0.5

OpenStudy (anonymous):

I had already gotten it right but thank you